OpenStudy (anonymous):

Solve x – 5 = 4 Solve x – 5 = 4 @Mathematics

jimthompson5910 (jim_thompson5910):

x – 5 = 4 x – 5 + 5 = 4 + 5 x + 0 = 9 x = 9

OpenStudy (anonymous):

x-5=4 +5 +5 x=9

OpenStudy (anonymous):

x=9